OpenStudy (anonymous):

HELP! The second of three numbers is 2 more than the first. The third number is twice the first. The sum of the first and third exceeds the second by 2. Find all three numbers

OpenStudy (cwrw238):

numbers can be written as x, x+2 and 2x

OpenStudy (cwrw238):

also x + 2x = x + 2 + 2

OpenStudy (anonymous):

now solve?

OpenStudy (cwrw238):

solve this equation for x yes

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

x = 2?

OpenStudy (cwrw238):

thats right so second number = 2 + 2 = 4 and 3rd = 2x = 4

OpenStudy (anonymous):

thanks so much!

OpenStudy (cwrw238):

yw